I'm having a hard time with this equation. Please help me!
y/8 = 4+ (y-2)/24

Answer by mananth(16946)

y/8 = 4+ (y-2)/24
y/8 -(y-2)/24 =4
LCM =24
3y-(y-2) /24 =4
3y-y+2 /24 =4
2y+2= 4*24
2(y+1)=96
y+2=48
y=48-1
y=47